# Cosmological Simulations of Stellar Halos with Gaia Sausage–Enceladus Analogues: Two Sausages, One Bun?

### v1

@@ Abstract
Observations of the Milky Way’s stellar halo find that it is predominantly comprised of a radially-biased population of stars, dubbed the Gaia Sausage–Enceladus, or GSE. These stars are thought to be debris from dwarf galaxy accretion early in the Milky Way’s history. Though typically considered to be from a single merger, it is possible that the GSE debris has multiple sources. To investigate this possibility, we use the IllustrisTNG50 simulation to identify stellar accretion histories in 98 Milky Way analogues—the largest sample for which such an identification has been performed—and find GSE-like debris in 32, with two-merger GSEs accounting for a third of these cases. Distinguishing single-merger GSEs from two-merger GSEs is difficult in common kinematic spaces, but differences are more evident through chemical abundances and star formation histories. This is because single-merger GSEs are typically accreted more recently than the galaxies in two-merger GSEs: the median infall times (with 16th and 84th percentiles) are $5.9_{-2.0}^{+3.3}$ and $10.7_{-3.7}^{+1.2}$ Gyr ago for these scenarios, respectively. The systematic shifts in abundances and ages which occur as a result suggest that efforts in modeling these aspects of the stellar halo prove ever-important in understanding its assembly.
Galactic archaeology (2178) — Stellar kinematics (1608) — Galaxy stellar halos (598)
@@
1 Introduction
One key prediction of the Lambda Cold Dark Matter ($\Lambda$CDM) cosmological paradigm is that of hierarchical structure formation (White & Rees, 1978; White & Frenk, 1991). In this model, the early Universe has a stochastically varying density distribution, in which overdensities serve as the seeds of galaxy formation. These overdensities interact gravitationally, growing by merging and accreting nearby matter. The assembly histories of galaxies like the Milky Way (MW) depend on the stochastic initial conditions of this density field, and each galaxy will form differently, even if the resulting halos are qualitatively similar. During the assembly process, infalling satellite galaxies are tidally disrupted by the strong gravitational forces of their host, leaving behind debris in “stellar halos” (Johnston et al., 1996; Johnston, 1998; Bullock & Johnston, 2005). The debris from these mergers can retain distinctive structure in chemodynamical space, thereby providing a useful approach to both identify and characterize these ancient events with present-day observations (Robertson et al., 2005; Font et al., 2006, 2011).
@@ 1 Introduction
The stellar halo is rich in structure, as summarized in Helmi (2020); it appears to be almost entirely comprised of debris from merging satellites (Deason et al., 2015, 2018; Haywood et al., 2018; Kruijssen et al., 2019; Naidu et al., 2020) and the in situ stars they heat out of the disk (Zolotov et al., 2009; Helmi et al., 2018; Belokurov et al., 2020, though see also Amarante et al. (2020)). The most significant accreted contribution to the inner halo is dubbed the Gaia Sausage–Enceladus, or GSE (Belokurov et al., 2018; Helmi et al., 2018)—see Deason & Belokurov (2024) for a recent review. The GSE stellar debris comprises more than 50% of the ex situ stars within 30 kpc of the Galactic center and has significantly radially-biased velocity anisotropy (Deason et al., 2018; Myeong et al., 2018; Lancaster et al., 2019; Necib et al., 2019a; Iorio & Belokurov, 2021; Naidu et al., 2021).
@@
Despite its well-studied chemodynamical properties, the underlying nature of the GSE debris is still being brought to light: many studies (Bonaca et al., 2020; Grunblatt et al., 2021; Feuillet et al., 2021; Hasselquist et al., 2021; Horta et al., 2024) find stars which are as young as $\sim 6$, rather than $\gtrsim 10$, Gyr old, but otherwise appear consistent with the GSE stellar population. One potential explanation for this is slow orbital decay allowing for ongoing star formation in the progenitor (Bonaca et al., 2020); however, another hypothesis is an additional dwarf galaxy component to the GSE debris. Rather than being from a single merger, these stars may have been contributed by multiple independent mergers. This hypothesis has been explored by Donlon et al. (2022), who argue that the structures present in chemodynamical spaces are more easily explained by multiple mergers than the single-merger GSE paradigm, a claim also consistent with a later infall time for GSE debris (Donlon et al., 2020, 2024).
Regardless of its origin, the ubiquity of the GSE debris in the inner halo makes it a useful testbed for assessing the consistency between expectations from hierarchical structure formation in the $\Lambda$CDM paradigm and observations of our Galaxy’s structure. In particular, models of galaxy formation must be consistent with GSE-like structures, in which a galaxy’s stellar halo is overwhelmingly composed of a population of radially-anisotropic stars, whether this population is from a single large merger or from multiple smaller mergers.
Assessing the prevalence of GSE-like debris in a cosmological context and discerning its origin requires an understanding of the range of possibilities allowed for the accretion history of MW-like galaxies in a CDM universe, which can be substantial due to cosmological variance. Fortunately, state-of-the-art large-volume, high-resolution simulations are expanding the possibility of quantifying the spread of expectations by providing substantial samples of MW-like galaxies that evolved in a cosmological setting. In this work, we use the IllustrisTNG project (Nelson, 2021), and the high-resolution TNG50 simulation (Nelson et al., 2019; Pillepich et al., 2019) in particular. [^1] Other comparable simulations include NewHorizon (Dubois et al., 2014) and FIREbox (Feldmann et al., 2023). See also Vogelsberger et al. (2020) for a recent review of cosmological simulations. Encompassing a volume of (50 comoving Mpc)3, TNG50 includes $\mathcal{O}(100)$ halos near the MW’s virial mass, resolving their subhalos down to nearly $10^{7}~{}\mathrm{M}_{\odot}$. This simulated sample of MW-like galaxies provides an unprecedented opportunity to characterize the spread in stellar properties in galaxies like our own, which by definition cannot be done with individual zoom-in simulations. Within this context, it is possible to better understand how common or rare the MW is among the range of possibilities. Moreover, it allows one to study large samples of GSE-like events, understanding how likely they are to form and from how many mergers.
Previous simulations have either focused on reproducing the properties of the GSE merger directly (e.g. Koppelman et al., 2020; Naidu et al., 2021; Amarante et al., 2022; Rey et al., 2023; Buch et al., 2024) or on selecting GSE-like events from a larger volume (e.g. Mackereth et al., 2019; Fattahi et al., 2019; Bignone et al., 2019; Elias et al., 2020; Grand et al., 2020; Evans et al., 2020; Dillamore et al., 2022; Belokurov et al., 2023; Rey et al., 2023; Orkney et al., 2023; García-Bethencourt et al., 2023). The majority of these works study $\mathcal{O}(10)$ MW analogues at most, limited in statistics by the scope of the simulation. Other large sample sizes exist, particularly in the Illustris simulation studied by Evans et al. (2020) and in the EAGLE simulation studied by Bignone et al. (2019), both of whom find $\mathcal{O}(100)$ MW analogues, but the present work provides a combination of large statistical power, state-of-the-art magnetohydrodynamics, and detailed phase-space criteria for GSE debris not present in previous studies.
In this work, we construct the assembly history of each MW-like galaxy in TNG50, identifying the mergers that contribute radially-anisotropic (RA) debris to its inner halo, as expected of the GSE. Roughly 30% of the TNG MWs host GSE-like events, although a smaller fraction (13% of MWs with GSE-like events, or 4% overall) match the extreme velocity anisotropy observed in our Galaxy’s GSE. We demonstrate that there is a reasonable probability that GSE-like debris is built from more than a single merger within a $\Lambda$CDM context and explore whether these merger components can be distinguished with kinematic, abundance, or age information.
This paper is organized as follows: §2 introduces TNG, the criteria used to select the sample of MW analogues, the procedure used to recover the merger history for each galaxy in the sample, and the tagging of individual star particles as belonging to a particular merger. §3 discusses the stellar halos across the sample and establishes a subsample of halos with RA stellar debris. §4 discusses the kinematic properties of this debris, and §5 discusses their star formation histories and chemical properties. Furthermore, §6 summarizes the primary conclusions of this work. Appendix A gives the IDs of MW analogues within the simulation, Appendix B provides additional technical details regarding the methodology (§2), and Appendix C contains additional figures for each MW analogue in the sample.
@@
*4
2 Methodology
2.1 Simulations
@@
The IllustrisTNG project (Nelson, 2021) is a suite of magnetohydrodynamic cosmological simulations performed with the moving-mesh code Arepo (Springel, 2010) and is a successor to Illustris (Vogelsberger et al., 2014a, b). [^2] All IllustrisTNG data, including post-processed catalogs, are available online at https://tng-project.org/. This work utilizes the highest-resolution simulation, TNG50 (Nelson et al., 2019; Pillepich et al., 2019, hereafter simply TNG), which covers a volume of $(51.7~{}\text{Mpc})^{3}$ from a redshift of $z=127$ to the present day according to cosmological parameters from the Planck Collaboration et al. (2016), viz. $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$. The simulation uses a dark matter particle mass resolution of $4.5\times 10^{5}~{}\mathrm{M}_{\odot}$ and a baryonic mass resolution of $8.5\times 10^{4}~{}\mathrm{M}_{\odot}$.
@@
Certain timesteps of the simulation are saved as “snapshots,” and within these snapshots, a halo-finding algorithm is used to track the growth and evolution of structure. This process is broken down to two steps. First, structures are identified with a friends-of-friends (FoF) algorithm (Davis et al., 1985), which is run only on the dark matter particles; other particle types (gas, stars, and black holes) are associated with the same FoF group as their nearest dark matter particle. These groups generally consist of a large central host halo and the subhalos in its potential well, though the FoF algorithm does not identify subhalos, nor does it require that the group be gravitationally bound. Next, bound structures (i.e., dark matter halos) are found within FoF groups using the Subfind algorithm (Springel et al., 2001; Dolag et al., 2009), which considers all particle types.
@@
There are a number of masses defined in the TNG FoF and Subfind catalogues relevant to this study. Each FoF group has a mass denoted $M_{200}$, which is the total mass within a sphere of radius $R_{200}$ drawn around the most bound particle of the FoF group such that the average density of that sphere is $200$ times the critical density of the universe. Since this is computed for FoF groups only, halos which are not the central objects in their group (i.e., satellites or other spatially-adjacent halos linked to the central halo by the FoF algorithm) do not have this quantity computed. However, each Subfind halo does have a total bound mass, hence denoted $M_{\mathrm{dyn}}$. [^3] We define a radius $R_{\mathrm{dyn}}$ from $M_{\mathrm{dyn}}$ such that a sphere with radius $R_{\mathrm{dyn}}$ and density 200 times the critical density has mass $M_{\mathrm{dyn}}$, the same relation between $R_{200}$ and $M_{200}$. This is the total mass of all particles bound to the halo, excluding the mass bound to any satellite halos. The contribution to $M_{\mathrm{dyn}}$ which comes from star particles is denoted $M_{\star}$. Since $M_{\mathrm{dyn}}$ and $M_{\star}$ are defined for all halos, not only those which are central in the FoF groups, these are the masses considered in this work.
@@
Because halos are identified in each individual snapshot, it is nontrivial to track the same physical object from snapshot to snapshot. To this end, the IllustrisTNG project provides a “merger tree,” which connects persistent physical structure between snapshots, recording the IDs given to the same structure at different points in time. Earlier snapshots hold “progenitors” of the later snapshots’ “descendant” halos. In TNG, these trees are found by the Sublink algorithm (Rodriguez-Gomez et al., 2015), though in this study, we modify the Sublink trees in cases where the algorithm fails to properly identify a descendant halo. More information about Sublink and our modifications is given in Appendix B.
2.2 Selection Criteria for Milky Way Analogues
Table 1: Selection Criteria for Milky Way Analogues
@@ 2.2 Selection Criteria for Milky Way Analogues
[…]
$>500$ kpc from halo with larger $M_{\mathrm{dyn}}$ 109
$>1$ Mpc from halo with $M_{\mathrm{dyn}}{}>10^{13}~{}\mathrm{M}_{\odot}$ 98
@@
A series of selection criteria are required to isolate galaxies within TNG that most closely resemble the MW. These analogues are chosen from the $z=0$ snapshot based on their stellar mass and two additional isolation criteria. In particular, they must
@@
The result of these selections is summarized in Table 1. Note that there are no particular requirements on the MW itself other than its stellar mass. In particular, there need not be a Large Magellanic Cloud (LMC) analogue, nor any particular stellar morphology.
The criterion on stellar mass is based on 68% confidence intervals quoted in the literature (Flynn et al., 2006; McMillan, 2011; Licquia & Newman, 2015; Bland-Hawthorn & Gerhard, 2016; Licquia & Newman, 2016; McMillan, 2017; Cautun et al., 2020, with Bland-Hawthorn & Gerhard (2016) and Licquia & Newman (2015) providing the widest upper and lower bound, respectively). This ensures that the selected galaxies at least qualitatively resemble the MW, though the range is intentionally left broad. There are a total of 131 galaxies in the TNG volume that satisfy this requirement.
@@
The second of the two isolation requirements is on a larger scale: the Local Group is known to be somewhat isolated, with no large galaxy groups in the nearby Universe (Karachentsev, 2005). To this end, the MW analogues are required to be more than 1 Mpc from any halo with total bound mass typical of massive groups, $M_{\mathrm{dyn}}\sim 10^{13}~{}\mathrm{M}_{\odot}$. 98 (90%) of the 109 galaxies which satisfy the first isolation requirement also satisfy this condition and these 98 halos comprise the final sample of MW analogues. Their TNG Subfind ID numbers are provided in Appendix A.
@@
The criteria used in this work are similar to, though distinct from, those chosen by Pillepich et al. (2023), who established a catalogue of MW and M31 analogues in the TNG volume. Though the majority (85%) of the sample defined above is included in their catalogue, there are three primary differences in selection criteria.
@@
First, this work focuses on MW analogues in particular, ignoring M31-like galaxies. Thus, the stellar mass criterion used here targets MW-mass galaxies alone. Furthermore, the stellar mass criterion pertains to the total stellar mass, not only the mass within 30 kpc; Pillepich et al. (2023) use the mass within 30 kpc to facilitate comparison to M31. All of the galaxies in our sample satisfy the Pillepich et al. (2023) stellar mass criterion, but the restricted range of $M_{\star}$ used in this work, which excludes M31, contributes in large part to the smaller sample size.
@@
Second, the Mpc-scale isolation criterion is performed on the level of Subfind halos rather than FoF groups in a way that is (for this sample) strictly looser. Pillepich et al. (2023) require that the MW resides in a FoF group whose host has an $M_{200}$ mass less than $10^{13}~{}\mathrm{M}_{\odot}$. However, we allow the MW analogue to be in a FoF group with a massive host so long as the analogue is sufficiently far ($>1$ Mpc) from the host. This allows for MW analogues that have been associated with a distant FoF group host to be included in the study while still ensuring the large-scale environment is empty of such large halos. Of the halos in our sample, four are excluded from Pillepich et al. (2023) due to their FoF group mass, and a further five are excluded because they are near a galaxy with $M_{\star}(<30~{}\mathrm{kpc})>10^{10.5}~{}\mathrm{M}_{\odot}{}$.
@@
Third, this study makes no requirements of the analogues’ stellar morphology, whereas Pillepich et al. (2023) require a stellar disk which is defined either in terms of the ellipticity of the stellar mass distribution or by visual inspection. The focus of this study is on accreted stellar populations, rather than the in situ stars which would comprise the majority of a disk component. While we have not checked this requirement directly, we infer that the MW analogues selected in this work do predominantly satisfy the morphology criterion: only six of our analogues are excluded from Pillepich et al. (2023) even though they satisfy the other criteria used in that work.
2.3 Recovering the Merger History

### v2

# Cosmological Simulations of Stellar Halos with Gaia Sausage–Enceladus Analogs: Two Sausages, One Bun?

@@ Abstract
Observations of the Milky Way’s stellar halo find that it is predominantly comprised of a radially biased population of stars, dubbed the Gaia Sausage–Enceladus, or GSE. These stars are thought to be debris from dwarf galaxy accretion early in the Milky Way’s history. Though typically considered to be from a single merger, it is possible that the GSE debris has multiple sources. To investigate this possibility, we use the TNG50 simulation to identify stellar accretion histories in 98 Milky Way analogs—the largest sample for which such an identification has been performed—and find GSE-like debris in 32, with two-merger GSEs accounting for a third of these cases. Distinguishing single-merger GSEs from two-merger GSEs is difficult in common kinematic spaces, but differences are more evident through chemical abundances and star formation histories. This is because single-merger GSEs are typically accreted more recently than the galaxies in two-merger GSEs: the median infall times (with 16th and 84th percentiles) are $5.9_{-2.0}^{+3.3}$ and $10.7_{-3.7}^{+1.2}$ Gyr ago for single- and two-merger scenarios, respectively. The systematic shifts in abundances and ages that occur as a result suggest that efforts in modeling these aspects of the stellar halo prove ever-important in understanding its assembly.
Unified Astronomy Thesaurus concepts: Galactic archaeology (2178); Stellar kinematics (1608); Galaxy stellar halos (598)
@@
[…]
The Astrophysical Journal, 983:119 (18pp), 2025 April 20
I Introduction
[…]
@@ I Introduction
The stellar halo is rich in structure, as summarized in Helmi (2020); it appears to be almost entirely comprised of debris from merging satellites (Deason et al., 2015, 2018; Haywood et al., 2018; Kruijssen et al., 2019; Naidu et al., 2020; Malhan et al., 2022; Malhan and Rix, 2024) and the in situ stars they heat out of the disk (Zolotov et al., 2009; Helmi et al., 2018; Belokurov et al., 2020, though see also Amarante et al. (2020)). The most significant accreted contribution to the inner halo is dubbed the Gaia Sausage–Enceladus, or GSE (Belokurov et al., 2018; Helmi et al., 2018)—see Deason and Belokurov (2024) for a recent review. The GSE stellar debris comprises more than 50% of the ex situ stars within 30 kpc of the Galactic center and has significantly radially-biased velocity anisotropy (Deason et al., 2018; Myeong et al., 2018; Lancaster et al., 2019; Necib et al., 2019a; Iorio and Belokurov, 2021; Naidu et al., 2021).
@@
Despite its well-studied chemodynamical properties, the underlying nature of the GSE debris is still being brought to light: many studies (Bonaca et al., 2020; Feuillet et al., 2021; Grunblatt et al., 2021; Hasselquist et al., 2021; Horta et al., 2024) find stars which are as young as $\sim 6$, rather than $\gtrsim 10$, Gyr old, but otherwise appear consistent with the GSE stellar population. One potential explanation for this is slow orbital decay allowing for ongoing star formation in the progenitor (Bonaca et al., 2020); however, another hypothesis is an additional dwarf galaxy component to the GSE debris. Rather than being from a single merger, these stars may have been contributed by multiple independent mergers. This hypothesis has been explored by Donlon et al. (2022), who argue that the structures present in chemodynamical spaces are more easily explained by multiple mergers than the single-merger GSE paradigm, a claim also consistent with a later infall time for GSE debris (Donlon et al., 2020, 2024).
@@
Assessing the prevalence of GSE-like debris in a cosmological context and discerning its origin requires an understanding of the range of possibilities allowed for the accretion history of MW-like galaxies in a CDM universe, which can be substantial due to cosmological variance. Fortunately, state-of-the-art large-volume, high-resolution simulations are expanding the possibility of quantifying the spread of expectations by providing substantial samples of MW-like galaxies that evolved in a cosmological setting. In this work, we use the IllustrisTNG project (Nelson, 2021), and the high-resolution TNG50 simulation (Nelson et al., 2019b; Pillepich et al., 2019) in particular. [^1] Other comparable simulations include NewHorizon (Dubois et al., 2014) and FIREbox (Feldmann et al., 2023). See also Vogelsberger et al. (2020) for a recent review of cosmological simulations. Encompassing a volume of (50 comoving Mpc)3, TNG50 includes $\mathcal{O}(100)$ halos near the MW’s virial mass, resolving their subhalos down to nearly $10^{7}~\mathrm{M}_{\odot}$. This simulated sample of MW-like galaxies provides an unprecedented opportunity to characterize the spread in stellar properties in galaxies like our own, leveraging the large volume of the simulation and correspondingly large sample size. Within this context, it is possible to better understand how common or rare the MW is among the range of possibilities. Moreover, it allows one to study large samples of GSE-like events, understanding how likely they are to form and from how many mergers.
Previous simulations have focused either on reproducing the properties of the GSE merger directly (e.g. Koppelman et al., 2020; Naidu et al., 2021; Amarante et al., 2022; Rey et al., 2023; Buch et al., 2024) or on selecting GSE-like events from a larger volume (e.g. Bignone et al., 2019; Fattahi et al., 2019; Mackereth et al., 2019; Elias et al., 2020; Evans et al., 2020; Grand et al., 2020; Dillamore et al., 2022; Belokurov et al., 2023; García-Bethencourt et al., 2023; Orkney et al., 2023; Rey et al., 2023). The majority of these works look for GSE analogs in $\mathcal{O}(10)$ MW analogs at most, limited in statistics by the scope of the simulation. Other works have studied GSE analogs with sample sizes comparable to (or greater than) our own, namely, in artemis (Dillamore et al., 2022), coco/colour (Bose et al., 2020), EAGLE (Bignone et al., 2019; Evans et al., 2020), and Illustris (Elias et al., 2020). Each of these works uses a sample of MW analogs comparable (or greater) in size to our own, but the present work provides a combination of large statistical power, state-of-the-art magnetohydrodynamics, and detailed phase-space criteria for GSE debris not present in previous studies.
[…]
This paper is organized as follows: Section II introduces TNG, the criteria used to select the sample of MW analogs, the procedure used to recover the merger history for each galaxy in the sample, and the tagging of individual star particles as belonging to a particular merger. Section III discusses the stellar halos across the sample and establishes a subsample of halos with RA stellar debris. Section IV discusses the kinematic properties of this debris, and Section V discusses their star formation histories (SFHs) and chemical properties. Furthermore, Section VI summarizes the primary conclusions of this work. A gives the IDs of MW analogs within the simulation, B provides additional technical details regarding the methodology (Section II), and C provides additional figures for each MW analog in the sample.
II Methodology
II.1 Simulations
@@
The IllustrisTNG project (Nelson et al., 2019b) is a suite of magnetohydrodynamic cosmological simulations performed with the moving-mesh code Arepo (Springel, 2010) and is a successor to Illustris (Vogelsberger et al., 2014a, b). [^2] All IllustrisTNG data, including post-processed catalogs, are available online at https://tng-project.org/. This work utilizes the highest-resolution simulation, TNG50 (Nelson et al., 2019a; Pillepich et al., 2019, hereafter simply TNG), which covers a volume of $(51.7~\text{Mpc})^{3}$ from a redshift of $z=127$ to the present day according to cosmological parameters from the Planck Collaboration et al. (2016), viz. $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$. The simulation uses a dark matter particle mass resolution of $4.5\times 10^{5}~\mathrm{M}_{\odot}$ and a baryonic mass resolution of $8.5\times 10^{4}~\mathrm{M}_{\odot}$.
@@
Certain time steps of the simulation are saved as “snapshots,” and within these snapshots, a halo-finding algorithm is used to track the growth and evolution of structure. This process is broken down into two steps. First, structures are identified with a friends-of-friends (FoF) algorithm (Davis et al., 1985), which is run only on the dark matter particles; other particle types (gas, stars, and black holes) are associated with the same FoF group as their nearest dark matter particle. These groups generally consist of a large central host halo and the subhalos in its potential well, though the FoF algorithm does not identify subhalos, nor does it require that the group be gravitationally bound. Next, bound structures (i.e., dark matter halos) are found within FoF groups using the Subfind algorithm (Springel et al., 2001; Dolag et al., 2009), which considers all particle types.
@@
There are a number of masses defined in the TNG FoF and Subfind catalogs relevant to this study. Each FoF group has a mass denoted $M_{200}$, which is the total mass within a sphere of radius $R_{200}$ drawn around the most bound particle of the FoF group such that the average density of that sphere is $200$ times the critical density of the Universe. Since this is computed for FoF groups only, halos that are not the central objects in their group (i.e., satellites or other spatially adjacent halos linked to the central halo by the FoF algorithm) do not have this quantity computed. However, each Subfind halo does have a total bound mass, hence denoted $M_{\mathrm{dyn}}$. [^3] We define a radius $R_{\mathrm{dyn}}$ from $M_{\mathrm{dyn}}$ such that a sphere with radius $R_{\mathrm{dyn}}$ and density 200 times the critical density has mass $M_{\mathrm{dyn}}$, the same relation between $R_{200}$ and $M_{200}$. This is the total mass of all particles bound to the halo, excluding the mass bound to any satellite halos. The contribution to $M_{\mathrm{dyn}}$ that comes from star particles is denoted $M_{\star}$. Since $M_{\mathrm{dyn}}$ and $M_{\star}$ are defined for all halos, not only those that are central in the FoF groups, these are the masses considered in this work.
@@
Because halos are identified in each individual snapshot, it is nontrivial to track the same physical object from snapshot to snapshot. To this end, the IllustrisTNG project provides a “merger tree,” which connects persistent physical structure between snapshots, recording the IDs given to the same structure at different points in time. Earlier snapshots hold “progenitors” of the later snapshots’ “descendant” halos. In TNG, these trees are found by the Sublink algorithm (Rodriguez-Gomez et al., 2015), though in this study, we modify the Sublink trees in cases where the algorithm fails to properly identify a descendant halo. More information about Sublink and our modifications is given in B.
II.2 Selection Criteria for MW Analogs
Table 1: Selection Criteria for Milky Way Analogs
@@ II.2 Selection Criteria for MW Analogs
$M_{\star}{}\in[4,\,7.3]\times 10^{10}~\mathrm{M}_{\odot}$ 131
>500>500 kpc from halo with larger $M_{\mathrm{dyn}}$ 109
>1>1 Mpc from halo with $M_{\mathrm{dyn}}{}>10^{13}~\mathrm{M}_{\odot}$ 98
@@
A series of selection criteria are required to isolate galaxies within TNG that most closely resemble the MW. These analogs are chosen from the $z=0$ snapshot based on their stellar mass and two additional isolation criteria. In particular, they must
@@
The result of these selections is summarized in Table 1. Note that there are no particular requirements on the MW itself other than its stellar mass; we do not select for any aspects of the Galaxy’s stellar morphology, even those that may be influenced by the assembly history. Further, there need not be a Large Magellanic Cloud (LMC) analog, an analog of M31, or other Local Group features. Such a local environment may be correlated with its formation history, and being agnostic to this may introduce systematic biases in the sample.
The criterion on stellar mass is based on 68% confidence intervals quoted in the literature (Flynn et al., 2006; McMillan, 2011; Licquia and Newman, 2015; Bland-Hawthorn and Gerhard, 2016; Licquia and Newman, 2016; McMillan, 2017; Cautun et al., 2020, with Bland-Hawthorn and Gerhard (2016) and Licquia and Newman (2015) providing the widest upper and lower bound, respectively). This ensures that the selected galaxies have a stellar mass comparable to the MW, though the range is intentionally left broad. There are a total of 131 galaxies in the TNG volume that satisfy this requirement.
@@
The second of the two isolation requirements is on a larger scale: the Local Group is known to be somewhat isolated, with no large galaxy groups in the nearby Universe (Karachentsev, 2005). To this end, the MW analogs are required to be more than 1 Mpc from any halo with a total bound mass typical of massive groups, $M_{\mathrm{dyn}}\sim 10^{13}~\mathrm{M}_{\odot}$. A total of 98 (90%) of the 109 galaxies that satisfy the first isolation requirement also satisfy this condition, and these 98 halos comprise the final sample of MW analogs. Their TNG Subfind ID numbers are provided in Appendix A.
It is important to note that the TNG box contains two Virgo-mass clusters ($M_{\mathrm{dyn}}\gtrsim 10^{14}~\mathrm{M}_{\odot}$) within the (51.7 Mpc)3 volume. The Virgo cluster itself is $\sim 16$ Mpc from the MW (Cantiello et al., 2024), but few of the MW analogs are this isolated: 79 (81%) of the 98 halos fall within 16 Mpc of one of the two largest clusters, and three analogs are close enough to belong to the FoF group of the largest cluster. Of the MW analogs that are found to host GSE-like debris (see Section III.2), the closest is 3.6 Mpc from one of the large clusters.
@@
The criteria used in this work are similar to, though distinct from, those chosen by Pillepich et al. (2023), who established a catalog of MW and M31 analogs in the TNG volume. Though the majority (85%) of the sample defined above is included in their catalog, there are three primary differences in selection criteria.
@@
First, this work focuses on MW analogs in particular, ignoring M31-like galaxies. Thus, the stellar mass criterion used here targets MW-mass galaxies alone. Furthermore, the stellar mass criterion pertains to the total stellar mass, not only the mass within 30 kpc; Pillepich et al. (2023) use the mass within 30 kpc to facilitate comparison to M31. All of the galaxies in our sample satisfy the Pillepich et al. (2023) stellar mass criterion, but the restricted range of $M_{\star}$ used in this work, which excludes M31, contributes in large part to the smaller sample size.
@@
Second, the Mpc-scale isolation criterion is performed on the level of Subfind halos rather than FoF groups in a way that is (for this sample) strictly looser. Pillepich et al. (2023) require that the MW resides in a FoF group whose host has an $M_{200}$ mass less than $10^{13}~\mathrm{M}_{\odot}$. However, we allow the MW analog to be in an FoF group with a massive host so long as the analog is sufficiently far (>1>1 Mpc) from the host. This allows for MW analogs that have been associated with a distant FoF group host to be included in the study while still ensuring the large-scale environment is empty of such large halos. Of the halos in our sample, four are excluded from Pillepich et al. (2023) due to their FoF group mass, and a further five are excluded because they are near a galaxy with $M_{\star}(<30~\mathrm{kpc})>10^{10.5}~\mathrm{M}_{\odot}{}$.
@@
Third, this study makes no requirements of the analogs’ stellar morphology, whereas Pillepich et al. (2023) require a stellar disk that is defined either in terms of the ellipticity of the stellar mass distribution or by visual inspection. The focus of this study is on accreted stellar populations, rather than the in situ stars that would comprise the majority of a disk component. While we have not checked this requirement directly, we infer that the MW analogs selected in this work do predominantly satisfy the morphology criterion: only six of our analogs are excluded from Pillepich et al. (2023) even though they satisfy the other criteria used in that work.
II.3 Recovering the Merger History